Progress Report

The new telescope was moved to the Devon Observatory and the first images were obtained on Saturday, April fifth,1997.
E mail from Dr. Hube :".....good images across the entire field with no obvious vignetting. The open cluster M67 in UBVR1 fits nicely within the 22' field of the camera and looks very pretty."

A number of minor problems still have to be worked out but good images are at last being obtained and Tyler Foster is now doing his honours undergraduate project on the system.

March,1999 update:

The Peltier Cooler in the CCD1 failed last year due to bad contacts or corrosion on a plug and socket assembly inside the camera. Our repairs in the lab solved the Peltier problem but probably should not have been undertaken outside of clean room/static grid facilities as a small anomaly on the chip was noticed afterwards.

The Thermo Electric Coolet (TEC) unit also had to be repaired after a number of chips overheated and blew. Apparently one of our astronomers didn't like the glow of the pilot light and threw his coat over the unit blocking the cooling air intake.

Last Summer the camera was returned to Spectra-Source for cleaning of the chip and the resolution of some other problems.

An interference pattern has shown up in some images. The pattern may be due to interference from the 57 khz. switching power supply that powers the Peltier device. I am currently looking for a 10 volt, 4 amp ,programmable constant current supply to kluge to the system and either verify or disprove the 'interference' theory.

Now if we could just do something about the cloudy weather of the last two months..

March to October update:

Construction of a new 'Eurocard' rack and pc boards for stepper motor drivers, etc. .

July, 1999:

Devon was shut down for the Summer and some components were returned to the labs for modification and repairs.

The filter changer center sensor had come adrift requiring disassembly of the filter rack. Notes on assembly were made in case of future trouble. The center sensor was re cemented with black epoxy before re-assembly and testing.

A Solution to the TEC Interference Problem

The CCD banding problem was traced to the TEC cycling the cooler on and off. The fix was to just set the TEC to a lower temperature causing the TEC power supply to run all the time. An external fan was added to cool the TEC control unit.

Tyler and Albert started work on better imaging software that also included focus, etc. controls.

September, 1999:

Tyler & Albert's dog & pony show at the observatory.

Everyone involved was invited out to the observatory for a demonstration of the telescope and Tyler and Albert's improved imaging software. My non-geeky wife amused herself by picking a gallon of the large 'Choke cherries' on the site.

The new imaging software is a vast improvement. There are still some problems with the marginal power of the filter change motor when the slide is near vertical. Warner Electric (Slo-Syn) has some high torque steppers and translators but they are expensive.

September, 2000:

Tyler brought the filter and camera assemblies in due to apparent failure of the focus control. Mice had nested in the rack power supply and caused some peripheral damage. Thorough cleaning, sterilizing, and repairs to some wiring were necessary.

I re-designed the filter control circuitry to incorporate a local oscillator that drives the filter stepper motor under program control and allows continuous position encoder readings. This system lifted vertically and held a 700 gram test weight using a visual basic control program. Tyler and Albert will now have to parse it into their "C" program but it may well save us the cost of a larger stepper motor.

Construction of a second experimental control rack was started using materials from the original system. The new (Rack2) system used an unregulated power supply with constant current drive to the focus and filter stepper motors. Camera position has been fixed in place as Tyler has told me not to bother with it for now.

October 2000

The filter change system (Rack1) that worked perfectly in the lab failed after being moved to Devon. Troubles were traced to complete failure of the filter position encoder. Current checks in the lab showed much lower than normal current draw by the encoder. I removed the cover in desperation and found that on this particular encoder the position sensor L.E.D. was acessible.

The failed L.E.D. in the Vernitech filter position encoder was replaced with an Osram LD271 which failed after being moved to the observatory. A limiting resistor was added to keep L.E.D. current below 20 ma with no apparent effect on output. A new BEI encoder is on order but this one works for now.

October 16th to 20th, 2000

Oscillator frequency is critical for the filter movement system, get it wrong and mechanical resonance renders the filter encoder useless. A screwdriver adjustable frequency trimpot was added to the rack system and tested OK in the lab.

The filter interlock switch had to be bypassed after the 0-80 screws that held the switch bracket in place sheared off.

January, 2001

Unseasonably warm weather and clear skies have allowed Tyler to do a series of observations and observe something formerly only observed by Radio Telescope.

March, 2001

The backup rack (Rack2) with current regulated focus and filter drives plus a new filter applet work very well on the 'Fake Telescope' mock up in the lab. Doug will arrange a trip to Devon for further testing sometime next week. Tyler assure me that the mice are gone.

BEI have sent the new encoder which has a different pinout than the old one but should fit mechanically. A pig tail converter plug has been made for testing during final installation on the telescope.

I have written help files and drawn schematics available in the parent directory.

Roy Schmaus March 6, 2001.

Back